OpenStudy (anonymous):

Jordan has 48 crayons. There are 10 crayons in each box. One box is missing 2 crayons. How many boxes does she have?

OpenStudy (anonymous):

5 boxes

OpenStudy (anonymous):

10*5=50-2=48